# Investigating Adversarial Vulnerability and Implicit Bias through Frequency Analysis

### v2

@@ 4.1 Data
For our experiments, we employed two benchmark image datasets: CIFAR-10 [21] and Imagenette [57]. CIFAR-10 consists of $60000$ RGB $32\times 32$ training images and $10000$ test images categorised into $10$ classes. When studying adversarial examples, we used the test images, and the training set was solely employed for fine-tuning models, as explained in greater detail in the subsequent section. Imagenette, a $10$-class subset of ImageNet [58] was used to scale up our experiments to a higher-dimensional data set. In this case, images were resized to $224\times 224$ and both training and test images (respectively, 9469 and 3925 images) were used to train modulatory masks, as detailed in Sec. 4.5.2.
4.2 Models
@@
To gain a more complete picture of how implicit bias and adversarial attacks are related in various scenarios, we employed classification models based on different neural architectures. For CIFAR-10 data, we used ResNet-20, a relatively small representative of the very well known ResNet family, introduced in [23]. We trained the model from scratch, and provide details on the training procedure in the Appendix (Sec. A.2). In the case of Imagenette, we employed both a CNN (ResNet-18 [23]) and a Vision Transformer (ViT-B/16 [59]). We took these models pretrained on ImageNet-1K and we only trained the final classification head, to adapt it to the classification task at hand.
4.3 Attacks
@@
We employed three $\ell_{\infty}$ gradient-based adversarial attacks: the Fast Minimum Norm (FMN) attack [22], Projected Gradient Descent (PGD) [13] and DeepFool [50]. All the attacks were employed in the untargeted setting. While FMN and DeepFool do not depend on any hyperparameter, for PGD it is necessary to allocate a perturbation budget $\epsilon$. For most experiments, we set it to the standard value of $\epsilon=\frac{8}{255}$ [60], but we provide an analysis of the robustness of our findings with respect to $\epsilon$ in Sec. 4.6. To implement all attack algorithms, we utilized the implementation provided in the Foolbox library [61, 62].
4.4 Mask training
@@
The key step in our experimental procedure is the training of Fourier masks (see Sec. 3.1). Starting from a trained, well-performing classifier, we freeze its parameters and prepend to it a pre-processing layer that computes the FFT of an image, multiplies it element-wise by the trainable mask and computes the inverse FFT. The real part of the resulting image is then fed into the classifier. The process of training the masks is identical for both the set of essential frequency masks and adversarial frequency masks, with the only difference being the data set used for mask training. We train essential frequency masks using clean images associated with their original labels. In contrast, for adversarial frequency masks, we utilize adversarial images and the adversarial labels produced by the classifier for those images. In this step, we optimize the standard Cross-Entropy loss function with the addition of an $\ell_{1}$ penalty term to promote mask sparsity. Further details on the mask training procedure are reported in Sec. A.3 in the Appendix.
4.5 Correlation between adversarial and essential frequency masks
4.5.1 CIFAR-10
@@
To provide evidence of the relation between the implicit bias of the network and the adversarial perturbations, we adopt a direct approach: we correlate the essential frequency masks with the adversarial frequency masks. This correlation analysis is performed using various methods, introduced in Sec. 3.2. Namely, we employ SVCCA [54] as a linear baseline, Cosine Similarity, and Intrinsic Dimension Correlation ($I_{d}$or [20]).
@@
Table 1: Correlation between essential frequency masks and adversarial frequency masks (CIFAR-10, ResNet-20). (SVCCA): SVCCA [54] correlation coefficient between essential and adversarial frequency masks; (Cosine sim) mean $\pm$ std. dev. cosine similarity between masks; ($\mathbf{I_{d}}$or) correlation coefficient between masks, according to $I_{d}$or [20]; ($\mathbf{P}$-value) P-value for $I_{d}$or (low means high probability of correlation)
@@
4.5.2 Imagenette
We now move to a larger-scale experiment using Imagenette [57], a 10-class subset of ImageNet [58]. The training of modulatory masks (essential frequency masks and adversarial frequency masks) was conducted following the same procedure used for CIFAR-10, with the only difference being that both the training and test images were used to calculate the masks. This choice was made due to the reduced size of Imagenette’s test set, which could impact the accuracy of subsequent analyses. Indeed an accurate estimation of intrinsic dimension is crucial when using $I_{d}$or [20] as the number of data points needed for reliable estimation scales exponentially with the intrinsic dimension itself [63]. For Imagenette, we compute the same correlation statistics as provided for CIFAR-10 (SVCCA, $I_{d}$or, and cosine similarities). However, due to the high resolution of masks (three channels of size $224\times 224$), repeated intrinsic dimension estimations are computationally infeasible on raw data. Therefore, we average-pool masks to $32\times 32$ pixels before applying correlation methods. The results are reported in Table 2. Similar to the previous dataset, the discrepancy between $I_{d}$or and SVCCA reveals the presence of a nonlinear correlation. Correlation is found for all combinations of adversarial attack (FMN, PGD, and DeepFool) and model architecture (ResNet and ViT).
Table 2: Correlation between essential frequency masks and adversarial frequency masks (Imagenette).
@@ 4.5.2 Imagenette
4.6 Correlation with variable perturbation magnitude
The PGD attack algorithm allows to tune the parameter $\epsilon$, which controls the magnitude of the adversarial perturbation in terms of $\ell_{\infty}$ norm. Throughout the paper, we kept this value fixed at $\frac{8}{255}$. However, varying this parameter can provide useful insight into our results in terms of correlation between essential frequency masks and adversarial frequency masks. Interestingly, as shown in Table 3, the correlation does not directly depend on the magnitude of the adversarial perturbation, and it emerges with high confidence even with more subtle attacks, with $\epsilon=\frac{1}{255}$ and $\epsilon=\frac{3}{255}$.
Table 3: Attack success rates and $I_{d}$or correlation results for PGD with variable perturbation magnitude $\epsilon$ on ResNet-20.
@@ 4.6 Correlation with variable perturbation magnitude
4.7 Class-specific content in masks
Expanding upon the findings presented in [19] regarding the clustering of modulatory masks, we propose the hypothesis that masks computed on images of the same class share substantial frequency content. To validate this hypothesis, we designed a simple test.
We sample multiple random subsets containing $k$ randomly chosen classes, with $k$ ranging from $1$ to $10$. Then, we apply $I_{d}$or to essential frequency and adversarial frequency masks belonging to that specific subset of classes. For each $k$, we consider the average correlation and P-value over the random subsets we sampled. As an example, for $k=1$ at each iteration we consider one random class alone, while for $k=10$ the subset of classes we are considering is not random at all, and we are correlating the masks for all images in the CIFAR-10 dataset.
If masks belonging to the same class shared common frequencies, we would anticipate the average correlation to increase and the average P-value to decrease (and, consequently, correlation confidence to increase) as we added more classes. This is because $I_{d}$or relies on a permutation test to compute the P-value and increasing the number of classes would lower the probability of matching masks belonging to the same class when they are permuted. In fact, in Fig. 4 a clear decreasing trend is visible for the correlation P-value versus $k$ (and, conversely, an increasing trend in the correlation coefficient). This trend is present with all the attacks we tested (FMN, PGD and DeepFool), indicating that there is a considerable amount of class-specific information encoded in the masks.
Figure 4: Average correlation according to $I_{d}$or between $k$-class sets of essential frequency and adversarial frequency masks (ResNet-20). Solid lines represent correlation coefficients, while faded dashed lines represent P-values.
Based on this observation, we envisioned the possibility of training a single mask that encodes the essential frequency content for an entire class. Such masks (one for each class, displayed in the Appendix A.5) can be obtained following the same approach used to learn essential frequency masks for single images, but training on all the images belonging to a certain class. We trained class-level masks on the training images of CIFAR-10 on ResNet-20 and observed that they effectively preserved correct classifications for the unseen test set, confirming that they capture a set of frequencies that covers those needed to classify each category. Motivated by these findings and by the strong correlation between adversarial attacks and essential frequencies, we move to evaluating whether essential frequency masks can help in the mitigation of adversarial attacks.
4.8 Modulatory masks can revert adversarial attacks
@@
At the current point, essential frequency masks are not sufficient to counter adversarial attacks. In fact, while it is true that they encode the frequency content required to predict a certain category, prior knowledge of the correct category is needed in order to choose the right mask to apply. However, a simple modification in the mask training strategy can address this issue.
@@
We retrain class-level essential frequency masks alternating the standard optimisation step, aimed at preserving correct classification of clean data, with a second step aimed at reverting adversarial attacks, akin to adversarial training [6]. Specifically, in this second step we consider a batch of adversarially perturbed images, all misclassified in a certain category, and optimise the mask, which we dub robust essential frequency mask, to revert the classification to the original category of each image.
@@
At test time, when an unseen and potentially adversarial sample is fed to the network, the prediction can be used to pick the corresponding mask, so that a second forward pass can be executed on the filtered image. This step proves helpful in counteracting adversarial attacks (Table 4), though it imposes a toll in terms of clean accuracy, as in the case of most adversarial defence techniques [51].
@@
Moreover, as reported in Table 4, in most cases robust essential frequency masks can generalise across different adversarial attack algorithms, meaning that exact knowledge of the type of threat one might be facing is not a crucial requirement. While this finding alone is insufficient to build a proper adversarial defence method, as an attacker may get access to the masks and produce new adversarial perturbations that deceive the masked model, we believe it represents a promising starting point for future research in this direction.
@@
Table 4: Clean and adversarial test accuracy when robust essential frequency masks are applied to ResNet-20 for CIFAR-10 classification. Rows report different adversarial attacks (including no attack, i.e. clean accuracy), while on the columns we report the results relative to masks trained on different attacks, including the base case when no mask is applied.
Data Mask
@@
5 Discussion
Our study investigates the relationship between adversarial attacks and the implicit bias of neural networks in the Fourier domain. We focus on standard network architectures like ResNets and ViTs and datasets such as CIFAR-10 and Imagenette, where direct mathematical analysis of implicit bias is particularly challenging.
For the first time, we computed the essential frequencies required for an adversarial image to be misclassified. We demonstrated that these frequencies are highly correlated with the essential frequencies needed for a clean image to be correctly classified, using multiple correlation metrics.
Our work represents an advancement in understanding the nature of adversarial attacks. Identifying the crucial frequencies utilised by attackers holds promising implications for developing new defence and detection algorithms, thereby enhancing the security and robustness of neural networks. Notably, by using a variant of the essential frequency masks to revert the effects of adversarially attacked images, we showed a new direction for generating adversarial defences.
@@ 5 Discussion
While studying the implicit bias of networks through the Fourier lens is a promising approach, that has been the focus of many recent works, it only offers a specific angle of analysis, which is a limitation of our study. However, our mask-based approach provides a flexible method for future investigations. For example, beyond examining the real part, researchers can investigate the essential Fourier phases in input images, opening new avenues for studying the implicit frequency bias of networks and their robustness. Additionally, other types of representations, such as wavelets, could be explored. Another limitation of our study is that it focuses on the input frequencies neglecting those of all the intermediate representations within the network. This area will be the subject of future investigations.
@@
In conclusion, our framework not only enhances the understanding of adversarial attacks in neural networks but also paves the way for developing more effective defence strategies, contributing to the overall advancement of AI security.
Appendix A Appendix
A.1 Computational Resources
In terms of computational requirements, our experimental procedure encountered two major resource-intensive tasks. First, training masks necessitates individual training for each image, which consumes substantial computational resources. Second, the computation of SVCCA and repeated intrinsic dimension estimates on large, high-dimensional mask data sets poses computational challenges. Both steps can benefit from hardware acceleration and were performed on a NVIDIA A100 GPU, equipped with 40GB of RAM.
A.2 Training details and hyperparameters
@@
As introduced in Sec. 4, we trained one model on CIFAR-10 [21] (ResNet-20 [23]) and two on Imagenette [57], namely a ResNet-18 [23] and a ViT-B/16 [59]. Here we provide additional details regarding the training (or fine-tuning) procedure of each model and the hyperparameters we chose.
@@
Our ResNet-20 implementation was taken from [64] and it follows the original architecture the authors of [23] proposed for CIFAR-10. This model was trained from randomly initialised weights for $200$ epochs, using Stochastic Gradient Descent with momentum, set to $0.9$. The learning rate was initially set to $0.1$, to be decayed by a factor $10$ twice, after $100$ and $150$ epochs. $\ell_{2}$ regularisation was employed, by means of a weight decay factor of $10^{-4}$. The final accuracy on the test set of CIFAR-10 was $92.29\%$.
@@
A.3 Details on the mask training procedure
For all the modulatory masks we produced, we employed Adam optimiser [65] with a learning rate of $0.01$. According to a criterion similar to early-stopping, masks were trained until convergence, and in any case for no less than $500$ optimisation steps each. We computed essential frequency masks only for correctly classified images, while adversarial frequency masks were trained exclusively for correctly classified images that were successfully made adversarial by the attack. Sparsity of masks was enforced by means of $\ell_{1}$ regularisation, weighted by a factor $\lambda=0.5$ for CIFAR-10 and $\lambda=0.01$ for Imagenette.
The same setup was employed for training the class-level masks introduced in Sec. 4.7 and 4.8, with the only difference that the training procedure was conducted on batches of images belonging to each class instead of individual images. In such case, the level of sparsity to be enforced on the masks is much lower, as they need to encode a set of frequencies that covers multiple images together. $\lambda$ was set to $0.01$ for essential frequency masks and to $0.001$ for robust essential frequency masks, as they require more fine-grained information to revert adversarial perturbations.
A.4 Examples of essential frequency masks
In Fig. 5 we provide some examples of essential frequency masks, computed on the same images employed to obtain the examples of adversarial frequency masks displayed in Fig. 3. Masks were trained starting from CIFAR-10 images on ResNet-20 architecture.
Figure 5: Examples of essential frequency masks.
A.5 Class-level essential frequency masks
Here we display the class-level masks we obtained on the training set of CIFAR-10, using the ResNet-20 model.
Figure 6: Class-level essential frequency masks (obtained on CIFAR-10 with ResNet-20).
As mentioned in the main text (Sec. 4.7), these masks successfully capture the essential frequency content needed to correctly classify test images, unseen during the training of the masks.
Moreover, we train class-level robust essential frequency masks (Sec. 4.8), that are able to revert the effect of the adversarial attack in most cases, restoring the correct label. When applied to clean data, such masks cause a slight drop in clean accuracy. Masks are trained with different sparsity requirements (regularisation $\lambda$ set to $0.01$ for essential frequency masks and to $0.001$ for robust essential frequency masks), to account for the high-frequency bias of adversarial attacks.
Figure 7: Class-level robust essential frequency masks (obtained on CIFAR-10 with ResNet-20, training on clean images and FMN-attacked images).

### v3

@@ IV-A Data
For our experiments, we employ two benchmark image datasets: CIFAR-10 [19] and Imagenette [54]. CIFAR-10 consists of $60000$ RGB $32\times 32$ training images and $10000$ test images categorised into $10$ classes. When studying adversarial examples, we use the test images, and the training set is solely employed for fine-tuning models, as explained in greater detail in the subsequent section. Imagenette, a $10$-class subset of ImageNet [55] is used to scale up our experiments to a higher-dimensional data set. In this case, we resize images to $224\times 224$ and use both training and test images (respectively, 9469 and 3925 images) to train modulatory maps, as detailed in Sec. IV-E2.
Figure 3: Examples of essential and adversarial frequency maps, represented as RGB images. The labels refer to the classification of the clean image. The maps were obtained using CIFAR-10 and the Fast Minimum Norm attack on ResNet-20.
IV-B Models
@@
To gain a more complete picture of how implicit bias and adversarial attacks are related in various scenarios, we employ classification models based on different neural architectures. For CIFAR-10 data, we use ResNet-20, a relatively small representative of the very well known ResNet family, introduced in [21]. We train the model from scratch, and provide details on the training procedure in the Appendix (Sec. A). In the case of Imagenette, we employ both a CNN (ResNet-18 [21]) and a Vision Transformer (ViT-B/16 [56]). We take these models pretrained on ImageNet-1K and we only train the final classification head, to adapt it to the classification task at hand.
IV-C Attacks
@@
We employ three $\ell_{\infty}$ gradient-based adversarial attacks: the Fast Minimum Norm (FMN) attack [20], Projected Gradient Descent (PGD) [12] and DeepFool [47]. All the attacks are considered in the untargeted setting. While FMN and DeepFool do not depend on any hyperparameter, for PGD it is necessary to allocate a perturbation budget $\epsilon$. For most experiments, we set it to the standard value of $\epsilon=\frac{8}{255}$ [57], but we provide an analysis of the robustness of our findings with respect to $\epsilon$ in Sec. IV-F. To implement all attack algorithms, we use the implementation provided in the Foolbox library [58, 59].
IV-D Map training
@@
The key step in our experimental procedure is the training of Fourier maps (see Sec. III-A). Starting from a trained, well-performing classifier, we freeze its parameters and prepend to it a pre-processing layer that computes the FFT of an image, multiplies it element-wise by the trainable map and computes the inverse FFT. The real part of the resulting image is then fed into the classifier. The process of training the maps is identical for both sets of modulatory maps, with the only difference being the data set used for map training. We train essential frequency maps using clean images associated with their original labels. In contrast, for adversarial frequency maps, we utilise adversarial images and the adversarial labels produced by the classifier for those images. In this step, we optimise the standard Cross-Entropy loss function with the addition of an $\ell_{1}$ penalty term to promote map sparsity. Further details on the map training procedure are reported in Sec. B in the Appendix.
IV-E Correlation between adversarial and essential frequency maps
IV-E1 CIFAR-10
@@
To provide evidence of the relation between the implicit bias of the network and the adversarial perturbations, we adopt a direct approach: we correlate the essential frequency maps with the adversarial frequency maps. This correlation analysis is performed using various methods, introduced in Sec. III-B. Namely, we employ SVCCA [51] as a linear baseline, Cosine Similarity, and Intrinsic Dimension Correlation ($I_{d}$or [18]).
@@
TABLE I: Correlation between essential frequency maps and adversarial frequency maps (CIFAR-10, ResNet-20). (SVCCA): SVCCA [51] correlation coefficient between essential and adversarial frequency maps; (Cosine sim) mean $\pm$ std. dev. cosine similarity between maps; ($\mathbf{I_{d}}$or) correlation coefficient between maps, according to $I_{d}$or [18]; ($\mathbf{P}$-value) P-value for $I_{d}$or (low means high probability of correlation)
@@
IV-E2 Imagenette
We now move to a larger-scale experiment using Imagenette [54], a 10-class subset of ImageNet [55]. The training of modulatory maps (essential frequency maps and adversarial frequency maps) is conducted following the same procedure used for CIFAR-10, with the only difference being that we use both the training and test images to calculate the maps. This choice is due to the reduced size of Imagenette’s test set, which could impact the accuracy of subsequent analyses. Indeed, an accurate estimation of intrinsic dimension is crucial when using $I_{d}$or [18] as the number of data points needed for reliable estimation scales exponentially with the intrinsic dimension itself [60]. For Imagenette, we compute the same correlation statistics as provided for CIFAR-10 (SVCCA, $I_{d}$or, and cosine similarities). However, due to the high resolution of maps (three channels of size $224\times 224$), repeated intrinsic dimension estimations are computationally infeasible on raw data. Therefore, we average-pool maps to $32\times 32$ pixels before applying correlation methods. The results are reported in Table II. Similar to the previous dataset, the discrepancy between $I_{d}$or and SVCCA reveals the presence of a nonlinear correlation. Correlation is found for all combinations of adversarial attack (FMN, PGD, and DeepFool) and model architecture (ResNet and ViT).
TABLE II: Correlation between essential frequency maps and adversarial frequency maps (Imagenette).
@@ IV-E2 Imagenette
IV-F Correlation with variable perturbation magnitude
The PGD attack algorithm allows to tune the parameter $\epsilon$, which controls the magnitude of the adversarial perturbation in terms of $\ell_{\infty}$ norm. Throughout the paper, we kept this value fixed at $\frac{8}{255}$. However, varying this parameter can provide useful insight into our results in terms of correlation between essential frequency maps and adversarial frequency maps. Interestingly, as shown in Table III, the correlation does not directly depend on the magnitude of the adversarial perturbation, and it emerges with high confidence even with more subtle attacks, with $\epsilon=\frac{1}{255}$ and $\epsilon=\frac{3}{255}$.
TABLE III: Attack success rates and $I_{d}$or correlation results for PGD with variable perturbation magnitude $\epsilon$ on ResNet-20.
@@ IV-F Correlation with variable perturbation magnitude
IV-G Modulatory maps can revert adversarial attacks
@@
We now move to assessing whether essential frequency maps can help mitigate the effect of adversarial attacks. The first step we take in this direction is overcoming the one-to-one relationship between input images and frequency maps by training a single map that encodes the essential frequency content for an entire class. Such class-level essential frequency maps can be obtained by following the same approach used to learn essential frequency maps for single images, but training on all the images belonging to a certain class. At this point, these maps would not suffice to counter adversarial attacks. In fact, while it is true that they encode the essential frequency content required to predict a certain category, prior knowledge of the correct category would be required in order to choose the right map to apply. However, a simple modification in the map training strategy can address this issue.
@@
Class-level essential frequency maps can be trained alternating the standard optimisation step, aimed at preserving correct classification of clean data, with a second step aimed at reverting adversarial attacks, akin to adversarial training [6]. Specifically, in the first step we consider a batch of clean images of category $c$ and aim for preserving the correct classification; in the second step we consider a batch of adversarially perturbed images, all misclassified in category $c$, and train the map to revert the classification to the original category of each image.
@@
At test time, when an unseen and potentially adversarial sample is fed to the network, the initial prediction can be used to pick the corresponding map, so that a second forward pass can be executed on the filtered image, to produce the final prediction. This step proves helpful in counteracting adversarial attacks (Table IV), though it imposes a toll in terms of clean accuracy, as in the case of most adversarial defence techniques [48].
@@
Moreover, as reported in Table IV, in most cases these adversarially-trained class-level essential frequency maps can generalise across different adversarial attack algorithms, meaning that exact knowledge of the type of threat one might be facing is not a crucial requirement. While these findings alone are insufficient to build an end-to-end adversarial defence method, as an attacker may get access to the maps and produce new adversarial perturbations that deceive the whole model, we believe they represent a promising starting point for future research in this direction.
@@
TABLE IV: Clean and adversarial test accuracy when class-level essential frequency maps are applied to ResNet-20 for CIFAR-10 classification. Rows report different adversarial attacks (including no attack, i.e. clean accuracy), while on the columns we report the results relative to maps trained on different attacks, including the base case where no map is applied.
Data Map
@@
V Discussion
Our study investigates the relationship between adversarial attacks and the implicit bias of neural networks in the Fourier domain. We focus on standard network architectures like ResNets and ViTs and datasets such as CIFAR-10 and Imagenette, where direct mathematical analysis of implicit bias is particularly challenging. For the first time, we compute the essential frequencies required for an adversarial image to be misclassified. We demonstrate that these frequencies are highly correlated with the essential frequencies needed for a clean image to be correctly classified, using multiple correlation metrics. Our work represents an advancement in understanding the nature of adversarial attacks and, more in general, the way the network represents the input. Apart from its impact on interpretability, identifying the crucial frequencies utilised by attackers holds promising implications for developing new defence and detection algorithms, thereby enhancing the security and robustness of neural networks. Notably, by using a variant of the essential frequency maps to revert the effects of adversarially attacked images, we show a new direction for crafting adversarial defences.
@@ V Discussion
While studying the implicit bias of networks through the Fourier lens is a promising approach, that has been the focus of many recent works, it only offers a specific angle of analysis, which is a limitation of our study. However, our map-based approach provides a flexible method for future investigations. For example, beyond examining the real part, researchers can investigate the essential Fourier phases in input images, opening new avenues for studying the implicit frequency bias of networks and their robustness. Additionally, other types of representations, such as wavelets, could be explored. Another limitation of our study is that it focuses on the input frequencies neglecting those of all the intermediate representations within the network. Moreover, our approach can be extended to analyse the correlation for a broader class of frequency representations, for instance coming from different networks or data distributions. These areas of research will be the subject of future investigations.
@@
In conclusion, our framework not only enhances the understanding of the way networks represent the input, with a particular focus on adversarially perturbed data, but it also paves the way for developing novel defence strategies, which could contribute to the overall advancement of AI security.
Acknowledgements
We acknowledge AREA Science Park for making the supercomputing platform ORFEO available for this work. Luca Bortolussi was supported by the PNRR project iNEST (Interconnected Nord-Est Innovation Ecosystem) funded by the European Union NextGenerationEU (Piano Nazionale di Ripresa e Resilienza (PNRR) – Missione 4 Componente 2, Investimento 1.5 – D.D. 1058 23/06/2022, ECS 00000043, CUP J43C22000320006).
Appendix A Training details and hyperparameters
@@
As introduced in Sec. IV, we trained one model on CIFAR-10 [19] (ResNet-20 [21]) and two on Imagenette [54], namely a ResNet-18 [21] and a ViT-B/16 [56]. Here we provide additional details regarding the training (or fine-tuning) procedure of each model and the hyperparameters we chose.
@@
ResNet-20 was trained from randomly initialised weights for $200$ epochs, using Stochastic Gradient Descent with momentum, set to $0.9$. The learning rate was initially set to $0.1$, to be decayed by a factor $10$ twice, after $100$ and $150$ epochs. $\ell_{2}$ regularisation was employed, by means of a weight decay factor of $10^{-4}$. The final accuracy on the test set of CIFAR-10 was $92.29\%$.
@@
Appendix B Details on the map training procedure
For all the modulatory maps we produced, we employed Adam optimiser [61] with a learning rate of $0.01$. According to a criterion similar to early-stopping, maps were trained until convergence, and in any case for no less than $500$ optimisation steps each. We computed essential frequency maps only for correctly classified images, while adversarial frequency maps were trained exclusively for correctly classified images that were successfully made adversarial by the attack. Sparsity of maps was enforced by means of $\ell_{1}$ regularisation, weighted by a factor $\lambda=0.5$ for CIFAR-10 and $\lambda=0.01$ for Imagenette.
The same setup was employed for training the class-level maps introduced in Sec. IV-G, with the only difference that the training procedure was conducted on batches of images classified in each category instead of individual images. In such case, the level of sparsity to be enforced on the maps is much lower, as they need to encode a set of frequencies that covers multiple images together. Hence, $\lambda$ was set to $0.001$.